OpenStudy (anonymous):

Limit Question:\[ \lim_{x\to0}\frac{x}{|x|} \]My hypothesis is that it doesn't exist.

OpenStudy (anonymous):

Problem is that l'hospital doesn't get me anywhere. \[ \frac{x'}{|x|'} = \frac{1}{\frac{x}{|x|}} = \frac{x}{|x|} \]

OpenStudy (anonymous):

One thing I have noticed is that: \[ \frac{x}{|x|} = \begin{cases} 1&\quad&x>0\\ -1&&x<0 \end{cases} \]

OpenStudy (anonymous):

So this means it can't exist because left and right hand don't exist.

OpenStudy (anonymous):

I mean the left and right hands exist, but don't converge

OpenStudy (jhannybean):

Is \(\ |x|\) even a one-to-one function? If thats even a valid question to ask....

OpenStudy (anonymous):

Well \(|x|\) is not one to one because \(|-1|=|1| = 1\), but I don't think that has any relevance when it comes to limits.

OpenStudy (anonymous):

I mean \(|x|\to 0\) as \(x\to0\), so it doesn't matter.

hartnn (hartnn):

Thats right, |x|/x or x/|x| doesn't exist as left and right hand limits are unequal.

hartnn (hartnn):

i mean their limit doesn't exist :P

OpenStudy (anonymous):

simply it does not exist 

OpenStudy (anonymous):

thus not differentiable at 0 :\

OpenStudy (anonymous):

I've got another question related to limits

OpenStudy (anonymous):

post